OpenStudy (anonymous):

PLEASE HELP , ILL GIVE A MEDAL A gardener wants to enclose a circular garden with a square fence, as shown below. If the circumference of the circular garden is about 48 feet, which of the following is the best estimate for the length of fencing needed? 61 feet 192 feet 31 feet 15 feet

OpenStudy (owlcoffee):

Well, if he wants to enclose the garden, wich is circular, we'd have to enclose it in a fencing wich is superior to the diameter of the circle. We know, that the circumference is calculated by: \[C _{\circ}= \pi (d)\] "d" is the diameter of the garden, so we know taht each side has to be longer or equal to it, so let's see what we know... We know that the circumference of the garden is 48ft, so let's replace it in the formula: \[48= \pi (d)\] and solving for "d" we would have found the vale of the diameter, wich in this case, is that the fence has to equal or surpass in length, let's see: \[\frac{ 48 }{ \pi }=d\] \[d=15.27\] And now, I'll let you choose the answer :)

OpenStudy (anonymous):

Thank you soooo much .

OpenStudy (anonymous):

@Owlcoffee that was wrong.....

OpenStudy (owlcoffee):

it was? I believe I misread?, I am very sorry for it :(

OpenStudy (owlcoffee):

Ah, I see, an "estimate" would be an estimation of 15.27, wich, because 2 is less than 5, we can call it 15. maybe?

OpenStudy (anonymous):

thats what i had chose , and it was wrong. but its okay i got the answer . thank you though

OpenStudy (anonymous):

So was the answer 61?

OpenStudy (anonymous):

nope , 192. i had to take the diameter and multiply it by 4.

OpenStudy (anonymous):

The diameter was calculated above to be 15.27. So 4 * 15.27= 61.12